If an 8kg mass is pulled up an incline at a constant speed, determine the tension in the string. And the coefficient of friction between the 8kg mass and the surface (Assume there is no friction between the cord and the pulley)


1
Expert's answer
2021-02-12T16:17:05-0500

"F=ma"


"T-\mu N-mg\sin\alpha=0"


"N=mg\cos\alpha"


"T-\mu mg\cos\alpha-mg\sin\alpha=0"


We obtained one equation with three unknowns. There is not enough data to solve the problem.